Tags:
create new tag
view all tags

20100408 Seed Laser Tuning & Locking

Today the seed laser would not lock to the external reference. The output power was checked and was ~450mW which was lower than expected (~500mW). Takano San tuned up the cavity by cleaning the mirrors and adjusting their alignment. He adjusted the SESAM and M3. See Seed Tuning for details on specific mirrors. The output power was increased to 520mW, but the laser still would not lock. Using the spectrum analyser, it was found that the reference signal was close to 359MHz as opposed to the laser (and ATF) frequency of 357MHz. 2MHz is far outside the laser's tuning range and so despite a higher output power, it couldn't lock.

After further investigation, the signal generator that synthesises the 357MHz frequency locally was found to not be operating properly. Changing the synthesis mode to "Auto" solved this.

The laser now locks stably.

Edit | Attach | Watch | Print version | History: r3 < r2 < r1 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r3 - 01 Jun 2010 - LaurieNevay

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback